# Compositional Flows for 3D Molecule and Synthesis Pathway Co-design

### v1

@@ 7 Discussion
Limitations. Our synthesis-based action space based on brick and linker synthons does not yet explore certain synthetic pathways including nonlinear synthesis pathway such as ring-forming reactions. This synthon-based generation is chosen to prevent the atoms from intermediate states being lost from reactions or forming new rings - which would degrade the accuracy of pose prediction for intermediate states. However, this constraint can be substantially mitigated by expanding the chemical search space through incorporating a larger building block library. For example, Sadybekov et al. (2021) successfully achieved a notable hit rate of 33% with brick-and-linker design.
Conclusion. In this work, we introduce Compositional Generative Flows (CGFlow), a flexible generative framework for jointly modeling compositional structures and continuous states. We introduce a simple extension to the flow matching interpolation process for handling compositional state transition. CGFlow enables the integration of GFlowNets for efficient exploration of compositional state spaces with flow matching for continuous state modeling. We apply CGFlow to 3D molecule and synthesis pathway co-design and develop 3DSynthFlow, which achieves state-of-the-art performance on both LIT-PCBA and CrossDocked benchmark. Future work includes using a more expressive model for pose prediction in 3DSynthFlow and developing more application-specific methods using CGFlow.
@@ D.2.1 State flow model
To model state flow for predicting ligand docking poses in 3DSynthFlow, we extend the Semla architecture introduced by Irwin et al. (2024). Semla is a scalable, $\mathrm{E}(3)$-equivariant model originally designed for 3D molecular generation; We refer reader to the original paper for full details. To adapt it for modeling 3D protein-ligand complexes, we incorporate protein encoding layers and protein-ligand message passing to capture interactions critical for binding, making the following modifications:
Let $P=\{1,\dots,M\}$ index protein residues. Specifically, we use GVP (Jing et al., 2021) to encode the protein pocket, producing an invariant feature vector for each residue ${\bm{h}}_{i}^{(\text{pro})}\in\mathbb{R}^{d},\;\forall i\in P$. The protein residue embeddings, along with their original positions ${\bm{x}}_{i}^{(\text{pro})}$, are combined with the ligand atom embeddings $({\bm{h}}_{j}^{(\text{lig})},{\bm{x}}_{j}^{(\text{lig})})$ for pairwise message passing at each modified Semla layer $(i,j)\in V\times P$, defined as:
@@
Since the atom type and bond connections are generated by the compositional flow model instead of the state flow model, therefore they remain fixed throughout the process. This means we cannot use techniques such as equivariant-OT to reduce the transport cost (Klein et al., 2023; Song et al., 2023), since they assume interchangeability between newly initialized nodes. We have also tried using harmonic prior from Jing et al. (2023); Stärk et al. (2024) to initialize position with bonding information as prior. We observed improved pose prediction with Gaussian noise during initial training, but no advantage of the harmonic prior over the Gaussian prior at convergence.
@@ D.2.2 Compositional flow model
To capture spatial relationships between the protein and ligand, the model uses a 3D protein-ligand complex graph for each state. The protein is represented as a residual graph, where each node corresponds to a $C_{\alpha}$ atom and connects to its K-nearest neighbors. Each ligand atom also connects to its K-nearest protein nodes. All pairwise distances between nodes are encoded in the graph edges, enabling the model to reason about the 3D structure. We represent each state $s$ as a molecular graph $s$ and include a GFlowNet condition vector $c$.
@@ E.1 Training details
Interpolation window $t_{window}$ 0.4
@@
Number of protein layers 6
@@
Time alpha $\alpha$ 1.0
Time beta $\beta$ 1.0
Time step $\Delta t$ 0.0125
@@
    Number of protein layers: the number of layers with protein-ligand message passing. This must be less or equal to the total number of layers.
@@
  3. 3.
    Time alpha: Defines the alpha parameter of the Beta distribution which time is sampled from.
@@
    Time beta: Defines the beta parameter of the Beta distribution which time is sampled from.
  5. 5.
    Time step: Defines the time step for inference.
@@
Learning rate ($P_{F}$, $P_{B}$) $10^{-4}$
State embedding dimension 128
@@
For action space subsampling, we randomly subsample 1% actions for AddFirstReactant and each bi-molecular reaction template $r\in\mathcal{R}_{2}$. However, for bi-molecular reactions with small possible reactant block sets $\mathcal{B}_{r}\in\mathcal{B}$, the memory benefit from the action space subsampling is small while a variance penalty is large. Therefore, we set the minimum subsampling size to 100 for each bi-molecular reaction, and the action space subsampling is not performed when the number of actions is smaller than 100.
@@ E.1.3 Training time and computational efficiency
The state flow model (i.e., the pocket-conditional pose predictor) was trained for 100 epochs with a batch size of 32 on 4 L40 GPUs (48GB), requiring a total of 18.4 hours. Since the state flow model is trained on the CrossDocked dataset and is reused across different test pockets, it incurs only a one-time computational cost. We plan to release the pretrained weights, so that users will need to train only the composition flow model tailored to their custom reward function and target.
In contrast, the composition flow model is trained individually for each pocket for 1,000 steps with a batch size of 64 and 80 flow matching steps. Training with GPU-accelerated docking takes between 12 and 20 hours (depending on the target) on a single A4000 GPU (16GB), making the computational requirement accessible for most practical drug discovery campaigns. Moreover, the composition flow model can be trained in a pocket-conditioned manner to enable zero-shot molecule sampling for any target pocket, thereby converting the training into a one-time cost in this setting.
@@
State Flow Model 4 $\times$ L40 (48GB) 32 100 epochs 18.4 hours
Composition Flow Model 1 $\times$ A4000 (16GB) 64 1,000 steps 12–20 hours
E.2 Datasets
E.2.1 LIT-PCBA Pockets
@@
E.2.2 CrossDocked2020
We train the State flow model of 3DSynthFlow on the commonly used CrossDocked dataset (Francoeur et al., 2020). We apply the splitting and processing protocol on the CrossDocked dataset to obtain the same training split of protein-ligand pairs as previous methods (Luo et al., 2021; Peng et al., 2022). To ensure fairness against baseline methods on the LIT-PCBA benchmark, we remove all PDB ids associated with LIT-PCBA proteins from the training set. Since we co-design 3D binding pose and synthesis pathway, unlike the pervious 2D-based methods, we can leverage this dataset for training the auxiliary State flow pose prediction model.
E.3 Baselines
@@ F.1 Reward computation without full docking
3DSynthFlow Full docking -11.97 ($\pm$ 0.07)
@@ F.3 Ablation study on time scheduling of 3DSynthFlow
3DSynthFlow ’s overlapping strategy, where positions are refined as synthons are added, clearly outperforms conventional autoregressive approaches - No overlap, adopted in previous works (See Appendix C).
@@ F.5 PoseCheck protein-ligand interactions
F.6 Synthesis success rate result on LIT-PCBA
Method Success Rate (%, $\uparrow$) Synthesis Steps ($\downarrow$)
FragGFN+SA 03.52 3.74
SynNet 47.50 3.45
BBAR 17.92 3.68
SynFlowNeta 54.60 2.55
SynFlowNetb 58.38 2.47
RGFN 47.43 2.46
RxnFlow 65.35 2.17
3DSynthFlow 53.12 3.22
Table. 14 shows that 3DSynthFlow attains comparable synthesizability to 2D baselines, demonstrating its generated molecules remain likely to be synthesizable. While the main contribution of our work is 3D co-design rather than improving synthesizability, these results confirm that our synthon-based 3D generation remains effective, with minimal trade-offs in synthetic feasibility.
F.7 Full sampling efficiency results
Table 15: Average number of diverse modes discovered versus the number of molecules explored. The best results are in bold.
@@ F.7 Full sampling efficiency results
3DSynthFlow 10.1 103.7 316.0 1607.0 4848.1
F.8 Full results for LIT-PCBA target
@@
3D Reaction 3DSynthFlow 3.34 ($\pm$ 0.07) 3.10 ($\pm$ 0.49) 3.45 ($\pm$ 0.12) 3.08 ($\pm$ 0.39) 3.06 ($\pm$ 0.62)
F.9 Example generation trajectories

### v2

@@ 7 Discussion
Limitations. Our synthesis-based action space based on arm and linker synthons does not yet explore certain nonlinear synthesis pathway such as ring-forming reactions. This synthon-based generation is chosen to prevent the leaving group atoms of intermediate states being lost or forming new rings – which would degrade the accuracy of pose prediction for intermediate states. However, this constraint can be substantially mitigated by expanding the chemical search space through incorporating a larger building block library. For example, Sadybekov et al. (2021) successfully achieved a notable hit rate of 33% with arm-and-linker design.
Conclusion. In this work, we introduce Compositional Generative Flows (CGFlow), a flexible generative framework for jointly modeling compositional structures and continuous states. We introduce a simple extension to the flow matching interpolation process for handling compositional state transition. CGFlow enables the integration of GFlowNets for efficient exploration of compositional state spaces with flow matching for continuous state modeling. We apply CGFlow to 3D molecule and synthesis pathway co-design and develop 3DSynthFlow, which achieves state-of-the-art performance on both LIT-PCBA and CrossDocked benchmark. Future work includes using a more expressive model for pose prediction and developing more application-specific methods using CGFlow.
@@
Acknowledgements
This work was supported by Basic Science Research Programs through the National Research Foundation of Korea (NRF), grant-funded by the Ministry of Science and ICT (RS-2023-00257479). This research was also supported by the NSERC Discovery Grant.
@@ D.2.1 State flow model
To model state flow for predicting ligand docking poses in 3DSynthFlow, we extend the Semla architecture introduced by Irwin et al. (2024). Semla is a scalable, $\mathrm{E}(3)$-equivariant model originally designed for 3D molecular generation; We refer reader to the original paper for full details for SemlaFlow. We make the following modifications to adapt Semla for 3D protein-ligand modelling:
Hierarchical pocket encoding: Inspired by AlphaFold3 (Abramson et al., 2024), which models protein at both atom level and token level, we propose a hierarchical encoding strategy HierSemla which first encodes each residues at atom level and aggregates atomic representation to residue level. More specifically, we introduce a Residue Encoder based on the Semla architecture, and apply it independently to encode each of its atom type, connections and position first. Then we perform mean pooling over the invariant and equivariant features within each residue, obtaining residue-level embedding $({\bm{h}}_{i}^{(\text{pro})},{\bm{x}}_{i}^{(\text{pro})})$. This hierarchical design enables all-atom resolution modeling of the protein pocket while significantly improving memory efficiency: by restricting full attention to within-residue interactions, we avoid the $\mathcal{O}(N^{2})$ memory cost of full atom-level attention across the entire protein, which can contain up to 2000 atoms.
Synthon attachment point information: In addition to the original atomic input features (atomic type and charge), we further provide the boolean feature is_attachment_point indicating whether this atom is the attachment point for a new synthon. Intuitively, this hints the model where the future synthon will be added in next reaction and bias the model to leave space for the future synthon in the intermediate state pose prediction. We find this additional feature lead to better pose prediction performance in practice.
Protein-ligand attention layer: The protein residue embeddings $({\bm{h}}_{i}^{(\text{pro})},{\bm{x}}_{i}^{(\text{pro})})$, along with their original positions ${\bm{x}}_{i}^{(\text{pro})}$, are combined with the ligand atom embeddings $({\bm{h}}_{j}^{(\text{lig})},{\bm{x}}_{j}^{(\text{lig})})$ for pairwise message passing at each modified Semla layer $(i,j)\in V\times P$, defined as:
@@
Modification to prior: Since the atom type and bond connections are generated by the compositional flow model instead of the state flow model, therefore they remain fixed throughout the process. This means we cannot use techniques such as equivariant-OT to reduce the transport cost (Klein et al., 2023; Song et al., 2023), since they assume interchangeability between newly initialized nodes. We have also tried using harmonic prior from Jing et al. (2023); Stärk et al. (2024) to initialize position with bonding information as prior. We observed improved pose prediction with Gaussian noise during initial training, but no advantage of the harmonic prior over the Gaussian prior at convergence.
@@ D.2.2 Compositional flow model
To capture spatial relationships between the protein and ligand, the model uses a 3D protein-ligand complex graph for each state. The protein is represented as a fully-connected residual graph, where each node corresponds to a $C_{\alpha}$ atom. The ligand is also connected to all protein nodes. All pairwise distances between nodes are encoded in the graph edges, enabling the model to reason about the 3D structure. We represent each state $s$ as an input of GFlowNet.
@@ E.1 Training details
Interpolation window $t_{window}$ 1.0
@@
Number of protein layers 4
Number of ligand layers 8
@@
Time distribution Beta(1.0, 1.0)
@@
    Number of protein layers: the number of layers with protein-protein message passing.
@@
    Number of ligand layers: the number of layers with protein-ligand message passing.
  3. 3.
@@
    Time distribution: Defines the time distribution
@@
Learning rate ($P_{F}$) $10^{-4}$
State embedding dimension 64
@@
For action space subsampling, we randomly subsample 1% actions for FirstSynthon and each bi-molecular reaction template $r\in\mathcal{R}_{2}$. However, for bi-molecular reactions with small possible reactant block sets $\mathcal{B}_{r}\in\mathcal{B}$, the memory benefit from the action space subsampling is small while a variance penalty is large. Therefore, we set the minimum subsampling size to 100 for each bi-molecular reaction, and the action space subsampling is not performed when the number of actions is smaller than 100.
@@ E.1.3 Training time and computational efficiency
The state flow model (i.e., the pocket-conditional pose predictor) was trained for 30 epochs on 8 A4000 GPUs (16GB), requiring a total of 26 hours. Since the state flow model is trained on the CrossDocked dataset and is reused across different test pockets, it incurs only a one-time computational cost. We plan to release the pretrained weights, so that users will need to train only the composition flow model tailored to their custom reward function and target.
In contrast, the composition flow model is trained for 1,000 iterations under 20 flow matching steps. Training with GPU-accelerated docking (UniDock; balance mode) takes between 7 and 12 hours (depending on the target) on a single A4000 GPU (16GB), making the computational requirement accessible for most practical drug discovery campaigns. Moreover, the composition flow model can be trained in a pocket-conditioned manner to enable zero-shot molecule sampling for any target pocket, thereby converting the training into a one-time cost in this setting.
@@
State Flow Model 8 $\times$ A4000 (16GB) dynamic 30 epochs 26 hours
Composition Flow Model 1 $\times$ A4000 (16GB) 64 1,000 steps 7–10 hours
E.2 Reward function
To train GFlowNet models, we employ the same reward function proposed in RxnFlow (Seo et al., 2024) and TacoGFN (Shen et al., 2024).
Pocket-specific optimization
. To optimize both QED andd Vina docking score, we set the reward function as:
$R(x)=w_{1}\text{QED}(x)+w_{2}\widehat{\text{Vina}}(x),$ (19)
where $\widehat{\text{Vina}}(x)$ is a normalized docking score. The parameters $w_{1}$ and $w_{2}$ serve as conditions of multi-objective GFlowNets (Jain et al., 2023b), and are set to 0.5 for non-GFlowNet baselines.
Pocket-conditional optimization
. We use the the modified reward function proposed by Shen et al. (2024). According to Seo et al. (2024), we remove SA score term (Ertl & Schuffenhauer, 2009) from the reward functions since 3DSynthFlow explicitly generate synthetic pathway:
$r_{\text{affinity}}(x)=\begin{cases}0&\text{if}~{}0\leq\text{Proxy}(x)\\ -0.04\times\text{Proxy}(x)&\text{if}~{}-8\leq\text{Proxy}(x)\leq 0\\ -0.2\times\text{Proxy}(x)-1.28&\text{if}~{}-13\leq\text{Proxy}(x)\leq-8\\ 1.32&\text{if}~{}\text{Proxy}(x)\leq-13\\ \end{cases}$
$r_{\text{QED}}(x)=\begin{cases}\text{QED}(x)/0.7&\text{if}~{}\text{QED}(x)\leq 0.7\\ 1&\text{otherwise}\\ \end{cases}$
$r_{\text{SA}}(x)=\begin{cases}\widehat{\text{SA}}(x)/0.8&\text{if}~{}\widehat{\text{SA}}(x)\leq 0.8\\ 1&\text{otherwise}\\ \end{cases}$
[…]
E.3 Datasets
E.3.1 LIT-PCBA Pockets
@@
E.3.2 CrossDocked2020
We train the State flow model of 3DSynthFlow on the commonly used CrossDocked dataset (Francoeur et al., 2020). We apply the splitting and processing protocol on the CrossDocked dataset to obtain the same training split of protein-ligand pairs as previous methods (Luo et al., 2021; Peng et al., 2022). We use 99,000 complexes as the training set and remaining 1,000 complexes as the validation set. Since we co-design 3D binding pose and synthesis pathway, unlike the pervious 2D-based methods, we can leverage this dataset for training the auxiliary State flow pose prediction model.
E.4 Baselines
@@ F.1 Reward computation without full docking
3DSynthFlow Full docking -11.82 ($\pm$ 0.07)
@@ F.3 Ablation study on time scheduling of 3DSynthFlow
  • Till end: all synthons are denoised jointly until $t=1$ ($\lambda=0.33,t_{window}=1.0$).
@@
We find 3DSynthFlow ’s Till end strategy, where all positions are refined until time $t=1$, clearly outperforms conventional autoregressive approaches - No overlap, adopted in previous works (See Appendix C). Our framework enables tuning of hyperparameter $\lambda$ and $t_{\text{window}}$ to accommodate varying data problems.
@@ F.5 PoseCheck protein-ligand interactions
F.6 Full sampling efficiency results
Table 16: Average number of diverse modes discovered versus the number of molecules explored. The best results are in bold. Results are computed over 2 seeds.
@@ F.6 Full sampling efficiency results
3DSynthFlow 9.6 71.2 242.4 1520.5 4432.3
F.7 Full results for LIT-PCBA target
@@
3D Reaction 3DSynthFlow 2.34 ($\pm$ 0.53) 2.20 ($\pm$ 0.41) 2.61 ($\pm$ 0.39) 2.73 ($\pm$ 0.47) 2.25 ($\pm$ 0.78)
F.8 Example generation trajectories